john can bake a cake in 10 minutes and ray can bake a cake and 7 minutes. how fast can they bake a cake together?
Answer by richwmiller(17219)

The cake needs more time to bake than 7 or 10 minutes.
x/7+x/10=1
10x+7x=70
17x=70
x=70/17
x=4.11764706